# Sampling as Bandits: Evaluation-Efficient Design for Black-Box Densities

### v1

@@ 3.3 Point-Selection Criterion
Proposition˜1 assures that, if the GP posterior converges, the GP-UJB also converges. Convergence analysis of the average-case GP regression error in the right-hand side is well established; see e.g. Kanagawa et al. (2018). The condition $g(\theta)<-1/2$ and $m_{n}(\theta)+k_{n}(\theta,\theta)<-1/2$, under which we have $C<1$, is mild when $\phi^{-1}=\log$. We can take a constant $0<B<0.5$ and assume $q(\theta)\leq B$ without loss of generality, because otherwise we can redefine the proportional $q(\theta)$ by dividing it by some constant so that the condition hold, where the density $p(\theta)$ remains invariant due to normalisation. By the bound, the log-proportional satisfies $g(\theta)=\log q(\theta)\leq\log B<-1/2$. Since $m_{n}(\theta)$ is the GP posterior mean that approximates $g(\theta)$ and $k_{n}(\theta,\theta)$ is bounded, the latter condition $m_{n}(\theta)+k_{n}(\theta,\theta)<-1/2$ holds by taking the bound $B$ small enough that $m_{n}(\theta)+k_{n}(\theta,\theta)$ does not exceed 1/2-1/2. Regardless, the latter condition is easy to verify numerically.
@@ 3.4 Hyperparameters of GP-UJB
Finally, we discuss the default setting of the hyperparameters of GP-UJB. The first hyperparameter of GP-UJB is the convex non-negative function $\phi$. The second hyperparameter of GP-UJB is a hyperparameter of the mean and covariance functions of the GP prior.
The balance between exploitation and exploration driven by GP-UJB is automatically adjusted by the choice of $\phi$. Choosing the function $\phi$ that results in a closed-form expression of GP-UJB is computationally appealing. Examples of closed-form GP-UJBs are summarised in Table˜1. Modelling the log-proportional $\log q(\theta)$ by GPs has been proven to be successful in multiple applications (e.g. Osborne et al., 2012; Järvenpää et al., 2021). Following these work, we suggest the exponential map $\phi(\cdot)=\exp(\cdot)$ as the default choice, which recovers the log-transformation $\phi^{-1}(\cdot)=\log(\cdot)$ as the inverse. This default choice is also supported by Remark˜2. Simulation studies on other choices of $\phi$ are provided in Section˜C.2.
Table 1: Examples of closed-form expressions of GP-UJB for different functions $\phi$, where we denote by $\varphi$ and $\Phi$ the probability density function and the cumulative distribution function of the standard normal distribution, respectively, and let $\sigma_{n}(\theta):=\sqrt{k_{n}(\theta,\theta)}$ for better presentation.
@@
$\max(0,x)$ $x$ $m_{n}(\theta)\varphi\left(m_{n}(\theta)/\sigma_{n}(\theta)\right)+\sigma_{n}(\theta)\Phi\left(m_{n}(\theta)/\sigma_{n}(\theta)\right)$
@@
Equipping the mean and covariance function of a GP prior with some hyperparameter $\psi$ is a common practice in GP regression. Finding an appropriate value of the hyperparameter $\psi$ maximises the performance of GP regression in general. Common examples of the hyperparameter $\psi$ include the length and variance constant of the Gaussian or Matérn covariance kernel. Typically, the hyperparameter $\psi$ is selected by maximising the log-likelihood of the GP posterior. For BO, using the maximum likelihood estimate of $\psi$ at each iteration is vital for a successful outcome. Following this standard practice, we use the maximum likelihood estimate of the hyperparameter $\psi$ for GP-UJB at each iteration.
@@
In sequential experimental designs for GPs, GP surrogate models at the first few iterations can be a poor approximation of a function of interest. This is because GPs can overfit to a very small number of data points. Selecting the first few points at random, without optimisation, is a widely-used strategy to improve stability of the GP surrogate model at early iterations. Algorithm 1 of BIS is immediately amenable for the same strategy, where for some integer $N_{0}>0$ the first $N_{0}$ points $\{\theta_{n}^{*}\}_{n=1}^{N_{0}}$ of the samples can be set to the first $N_{0}$ points of the scaled Halton sequence without optimisation.
Remark 4 (Limitations).
In this work, we limit our focus to densities in low-dimensional domains. Difficulties in both scalability and accuracy of GPs in high dimensions is a well-recognised open challenge. Besides, sampling from densities in high dimensions is an active research field of Monte Carlo methods in general. Assuming both high dimensionality and computational expensiveness of the target density incurs two nested challenges in the problem setup. We aim to solve one problem at a time, focusing on the case where the target density is computationally expensive but in low dimensions. Still, this case has numerous important applications. For example, the Lorenz weather forecast model depends on a 40-dimensional differential equation in the data domain, while the parameter space is two-dimensional (Wilks, 2005). Another example is a hierarchical model for very large satellite spatial data (Lyne et al., 2015), where the majority of the parameters can be analytically marginalised out, yet a few remaining parameters require sampling.
4 Convergence Analysis
This section establishes the weak convergence of the weighted samples $\{\theta_{n}^{*},w_{n}^{*}\}_{n=1}^{N}$ obtained by Algorithm 1. Firstly, Section˜4.1 introduces the definition of maximum mean discrepancy (MMD; Smola et al., 2007; Gretton et al., 2012), which we use as a metric to measure the approximation error between the weighted samples and the target density. Section˜4.2 then derives the convergence rate of Algorithm 1 in our default setting where the proposal density is uniform. Section˜4.3 provides the convergence rate of Algorithm 1 in a general setting where the proposal density is arbitrary.
4.1 Metric of Convergence
Consider a metric in the space of probability distributions on $\Theta$, which metrises the weak convergence of probability distributions. That is, a probability distribution weakly converges to another if and only if the metric between them converges to zero. A convenient approach to proving the convergence between the BIS samples and the target density $p$ is to show the decay of the metric between them. In this regard, MMD is a computationally and theoretically appealing choice of metric.
MMD is a metric determined by a kernel function $\kappa:\Theta\times\Theta\to\mathbb{R}$. Recall that, for every kernel, there exists a uniquely associated Hilbert space of functions on $\Theta$, called reproducing kernel Hilbert space (RKHS). See e.g. Paulsen and Raghupathi (2016) for introduction to RKHS. Denote by $\mathcal{H}_{\kappa}$ the RKHS of the kernel $\kappa$, with the norm denoted $\|\cdot\|_{\mathcal{H}_{\kappa}}$. MMD is then defined as the worst-case error between two expectations with respect to $p$ and $q$:
MMD satisfies all the axioms of a metric and metrises the weak convergence of probability distributions under mild regularity of the kernel (Simon-Gabriel et al., 2023). For our theoretical analysis, the choice of kernel $\kappa$ is not particularly crucial. Our convergence bounds do not depend on the choice of kernel $\kappa$ up to a condition posited below. If there exists any kernel that satisfies the condition, it suffices for our analysis on the weak convergence of the BIS samples.
A kernel $\kappa$ is said to be integrally strictly positive definite if $\int_{\Theta\times\Theta}\kappa(\theta,\theta^{\prime})d\mu(\theta)d\mu(\theta^{\prime})>0$ for all finite non-zero signed Borel measures $\mu$ on $\Theta$. This is analogue to strictly positive definiteness of kernels. For example, Gaussian kernels satisfy this property (Sriperumbudur et al., 2010). For convenience in our analysis, we posit the following condition on the kernel $\kappa$ used for MMD. In the condition, all the bounds of the kernel are set to $1$ with no loss of generality. This is because, if the bounds of the kernel exceeds $1$, we can multiply the kernel by some constant and define a new kernel whose bounds are below $1$. Gaussian kernels are a trivial example that meets the condition.
Kernel Condition: The kernel $\kappa:\Theta\times\Theta\to\mathbb{R}$ is integrally strictly positive definite. The kernel $\kappa$ admits a factorisation $\kappa(\theta,\theta^{\prime})=\prod_{i=1}^{d}\kappa_{i}(\theta_{(i)},\theta_{(i)}^{\prime})$ for some kernels $\kappa_{1},\dots,\kappa_{d}$, where $\theta_{(i)}$ and $\theta_{(i)}^{\prime}$ denote the $i$-th coordinate of $\theta$ and $\theta^{\prime}$. For each $i=1,\dots,d$, the kernel $\kappa_{i}$ is uniformly bounded by $1$ and admits the derivative $(\partial/\partial\theta_{(i)})(\partial/\partial\theta^{\prime}_{(i)})\kappa(\theta_{(i)},\theta^{\prime}_{(i)})$ uniformly bounded by $1$.
4.2 Deterministic Rate in Uniform Case
Our first analysis focuses on the case where the proposal density $u$ is uniform, as discussed in Section˜3.2. Strikingly, the uniformity of $u$ results in a deterministic convergence rate, regardless of whether the proposal sequence is random or deterministic. A key quantity in this analysis is the star discrepancy $\operatorname{D}^{*}$ of given $K$ points $\{\theta_{n}\}_{n=1}^{K}$ in $\Theta$, which we define by
$\operatorname{D}^{*}\left(\{\theta_{n}\}_{n=1}^{K}\right):=\sup_{B\in\mathcal{B}}\left|\frac{\#\{\theta_{n}\in B\}}{K}-\frac{\operatorname{Vol}(B)}{\operatorname{Vol}(\Theta)}\right|$ (9)
where $\mathcal{B}$ is a set of all subboxes in $\Theta$ s.t. the left-boundary at every i-th coordinate is equal to that of $\Theta$, $\{\theta_{n}\in B\}$ denotes the number of the points contained in the box $B$, and $\operatorname{Vol}(\cdot)$ denotes the volume of a given region. Although the star discrepancy is typically defined for points in the unit cube, we use the above definition generalised to the domain $\Theta$.
Denote by $\partial_{i}f(\theta)$ the partial derivative of a function $f(\theta)$ with respect to the i-th coordinate of the argument $\theta$. Denote by $\partial_{1:d}f(\theta)$ the first mixed partial derivative of a function $f(\theta)$, that is, $\partial_{1:d}f(\theta)=\partial_{1}\cdots\partial_{d}f(\theta)$. In what follows, denote by $\delta_{N}$ the empirical distribution of the $N$ weighted samples $\{\theta_{n}^{*},w_{n}^{*}\}_{n=1}^{N}$ of BIS. We arrive at the main result.
Theorem 1.
Suppose that the target density $p$ admits the first mixed partial derivative $\partial_{1:d}p(\theta)$ that is uniformly continuous. Suppose that the proposal density $u$ is uniform. Given any proposal sequence $\{\theta_{n}\}_{n=1}^{M+N}$ and criterion $U_{n}$, the weighted samples $\{\theta_{n}^{*},w_{n}^{*}\}_{n=1}^{N}$ of Algorithm 1 achieve
$d(p,\delta_{N})\leq C_{1}\operatorname{D}^{*}\left(\{\theta_{n}\}_{n=1}^{N+M}\right)+C_{2}\frac{M}{N+M}$ (10)
where $C_{1}$ and $C_{2}$ are constants dependent only on $p$ and $u$.
Proof.
The proof is provided in Section˜B.2. ∎
Theorem˜1 provides an immediate corollary on an explicit convergence rate of BIS. It suffices to know a choice of the proposal sequence $\{\theta_{n}\}_{n=1}^{M+N}$ that gives an upper bound of the star discrepancy. The Halton sequence leads to a well-known bound of the star discrepancy (see e.g. Niederreiter, 1992). We obtain the following rate under our default setting in Section˜3.2.
Corollary 1.
In the setting of Theorem˜1, suppose that the proposal sequence $\{\theta_{n}\}_{n=1}^{M+N}$ is the scaled Halton sequence in Definition˜1. The weighted samples $\{\theta_{n}^{*},w_{n}^{*}\}_{n=1}^{N}$ of Algorithm 1 achieve
$d(p,\delta_{N})\leq C_{1}\frac{\log(N+M)^{d}}{N+M}+C_{2}\frac{M}{N+M}$ (11)
where $C_{1}$ and $C_{2}$ are constants dependent only on $p$ and $u$.
Proof.
The proof is provided in Section˜B.3. ∎
This rate is deterministic and faster than the concentration rate of standard Monte Carlo integration $\mathcal{O}(N^{-1/2})$. Interestingly, even i.i.d random samples from the uniform density $u$ lead to another bound of the star discrepancy, which holds almost surely (Morokoff and Caflisch, 1994). We do not consider this case, as the convergence rate in this case is slower. Still, it is worth noting that the i.i.d random proposal sequence can result in the almost sure convergence of the MMD. Finally, the above convergence rate implies the weak convergence of the weighted samples.
Corollary 2.
Suppose the same setting as Corollary˜1. The empirical distribution $\delta_{N}$ of the weighted samples $\{\theta_{n}^{*},w_{n}^{*}\}_{n=1}^{N}$ weakly converges to the target density $p$ as $N\to\infty$.
Proof.
The proof is provided in Section˜B.6. ∎
Notably, the convergence rate is independent of a choice of the criterion $U_{n}$. It guarantees that BIS will not worsen the convergence rate of importance sampling. Therefore, a reasonably designed point-selection step in BIS will only accelerate the convergence.
4.3 Concentration Rate in General Case
Our next analysis considers a general setting, where the proposal density $u$ is arbitrary and the proposal sequence $\{\theta_{i}\}_{i=1}^{\infty}$ is a sequence of i.i.d. random samples from $u$. Since the proposal sequence $\{\theta_{i}\}_{i=1}^{\infty}$ is stochastic, we derive a concentration rate of the MMD between the weighted samples and the target density.
Theorem 2.
Suppose that the proposal density $u$ is a density uniformly bounded on $\Theta$. Given i.i.d. random samples from $u$ as a proposal sequence $\{\theta_{n}\}_{n=1}^{M+N}$ and a criterion $U_{n}$, the weighted samples $\{\theta_{n}^{*},w_{n}^{*}\}_{n=1}^{N}$ of Algorithm 1 achieve that, for all $0<\epsilon<1$,
where $\mathbb{P}$ is a probability taken with respect to the random samples $\theta_{1},\dots,\theta_{M+N}\overset{i.i.d.}{\sim}u$, and $C_{1}$, $C_{2}$, and $C_{3}$ are constants dependent only on $p$ and $u$.
Proof.
The proof is provided in Section˜B.5. ∎
We have an immediate corollary on the almost-sure weak convergence of the weighted samples.
Corollary 3.
Suppose the same setting as Theorem˜2. The empirical distribution $\delta_{N}$ of the weighted samples $\{\theta_{n}^{*},w_{n}^{*}\}_{n=1}^{N}$ weakly converges to the target density $p$ a.s. as $N\to\infty$.
Proof.
The proof is provided in Section˜B.6. ∎
While this setting allows for any proposal density $u$, the resulting convergence rate is stochastic and slower than the rate established in Section˜4.2. This result also motivates our default setting of the proposal density and the proposal sequence discussed in Section˜3.2. This completes the convergence analysis of BIS. Next we provide empirical assessment of the performance of BIS.
@@ 5 Empirical Assessment
This section examines the empirical performance of BIS in three distinct experiments. The first experiment concerns three benchmark densities in two dimensions, where we illustrate the samples obtained by BIS, together with the convergence rate. The second experiment considers a weather forecast model from Lorenz (1995), for which Bayesian inference is performed using a synthetic likelihood based on simulations (Hakkarainen et al., 2012). The third experiment focuses on exact Bayesian inference of the g-and-k models (Prangle, 2020), whose density requires a numerical optimisation to be evaluated at each data point. The source codes of the experiments are available in https://github.com/takuomatsubara/Bandit-Importance-Sampling.
@@ 5.1 Benchmark Densities
We consider three benchmark densities. The first density is a simple unimodal Gaussian density. The second density is a bimodal exponential-family density illustrating a challenging scenario that involves multiple modes and benefits from efficient exploration of the domain. The last density is a banana-shaped density illustrating another challenging scenario that involves a complex tail probability. All the densities are defined in the following form proposed in Järvenpää et al. (2021):
$p(\theta)\propto\exp\bigg{(}-\frac{1}{2}\begin{bmatrix}T_{1}(\theta)\\ T_{2}(\theta)\end{bmatrix}^{\mathrm{T}}\begin{bmatrix}1&\rho\\ \rho&1\end{bmatrix}\begin{bmatrix}T_{1}(\theta)\\ T_{2}(\theta)\end{bmatrix}\bigg{)}$, (13)
where Section˜D.1 contains the definition of $T_{1}$, $T_{2}$, and $\rho$ for each density. The domain $\Theta$ of each density is, respectively, $\Theta=[-16,16]^{2}$, $\Theta=[-6,6]^{2}$, and $\Theta=[-6,6]\times[-20,2]$. The aim of this experiment is to demonstrate the performance of BIS under each benchmark scenario.
@@
Figure 2: Visualisation of 100 BIS samples obtained for each benchmark density. The top panels show the contour colourmap of each density. The bottom panels show the initial 10 random points (star) and the following 90 points selected by BIS (circle). In the bottom panels, we added the contour plot of each density (solid line), where the contour values of each density were powered to $1/3$ for better visualisation of the curvature of the tail geometry.
By BIS, we obtained 100 weighted samples that approximate each target density. Following Järvenpää et al. (2021), the first 10 points were randomly selected from the Halton sequence, in order to increase stability of the GP surrogate model. The mean and covariance function of the GP prior in GP-UJB were set to, respectively, zero and the Gaussian kernel with its length-scale and variance constants as hyperparameters to tune. The size of the candidate pool $M$ in BIS was fixed to $M=2048$ for all the densities. See Section˜C.3 for simulation studies on sensitivity of BIS to the size of the candidate pool $M$. Figure˜2 compares each target density and the obtained samples, demonstrating that BIS captures the high probability region of each density. In particular, BIS efficiently tracked down the complex geometry of the bimodal and banana-shaped densities.
We measured the approximation error of the obtained samples by MMD. For computation of the MMD, we approximated each ground-truth target density by standard importance sampling based on 100,000 points from the scaled Halton sequence. We used the 100,000 weighted samples as a sufficiently accurate representation of each target density. Figure˜3 shows the MMD between the target density and the BIS samples. For reference, BIS was compared with standard importance sampling, where we obtained samples from the scaled Halton sequence and assigned self-normalised weights to them. Section˜D.1 further shows the number of samples required for standard importance sampling to surpass the approximation error of 100 BIS samples. BIS reduced the number of samples required to achieve the approximation error by nearly 95% on average.
@@
Figure 3: The approximation error of BIS weighted samples (solid line) and importance sampling based on the QMC sequence (dotted line). The experiment was repeated 10 times, where the bold line represents the averaged error and the band represents the 95% confidence interval.
BIS does not require the GP surrogate model to converge exactly to the target density, where a reasonable approximation of the target density suffices for guiding sample selection. That said, it may still be instructive to measure the approximation error between the target density and the GP surrogate model used in BIS. We computed the total variation distance (TVD) between the target density and the plug-in density estimator based on the GP mean $m_{n}$. Figure˜4 demonstrates the decay of the approximation error. The plug-in density estimator of BIS was compared with three other approaches of GP-based density estimators. The first approach is a naive plug-in density estimator based on the GP mean trained on QMC samples. The second approach is a modification of Sinsbeck and Nowak (2017) derived in Järvenpää et al. (2021), called the expected integrated variance (EIV) design. The EIV design is numerically more stable than the original proposal in Sinsbeck and Nowak (2017) since it needs no simulation from the GP posterior. The last approach is the randomised BO of Kim and Sanz-Alonso (2024). Note that the mean and covariance function of the GP surrogate model, including the hyperparameter, is the same in each method, where the only difference is in the input points on which the GP is trained. The GP surrogate model trained on the weighted samples of BIS achieves the least error in all the cases.
@@
Figure 4: The approximation error of the GP-based density estimator trained on BIS samples (solid line), the QMC sequence (dotted line), points by the randomised BO (dashed line), and points by the EIV design (dash-dotted line). The experiment was repeated 10 times, where the bold line represents the averaged error and the band represents the 95% confidence interval.

### v2

@@ 5.2 Approximation Error of GP-UJB
The last conditions $g(\theta)<-1/2$ and $m_{n}(\theta)+k_{n}(\theta,\theta)<-1/2$ are mild when employing the log-transformation $\phi^{-1}(\cdot)=\log(\cdot)$. Crucially, the target density $p(\theta)$ is invariant to the scaling of the proportional $q(\theta)$ due to normalization. Consequently, $q(\theta)$ can be arbitrarily rescaled by a constant factor without altering the original density $p(\theta)$. Without loss of generality, we may therefore assume that $q(\theta)$ is bounded by a constant $B$, s.t., $g(\theta)=\log q(\theta)\leq\log B<-1/2$. By similar reasoning, a sufficiently small scaling factor ensures that the GP approximation also satisfies that $m_{n}(\theta)+k_{n}(\theta,\theta)<-1/2$. Furthermore, this latter condition is readily verifiable numerically.
The proof of Proposition˜1 is provided in Supplement˜B. Proposition˜1 assures that, if the GP posterior converges, the GP-UJB also converges. Convergence analysis of the average-case GP regression error is well established; see, e.g., (Kanagawa2018).
@@ 5.3 Hyperparameters of GP-UJB
Finally, we discuss the specification of the hyperparameters for GP-UJB: the convex non-negative function $\phi$ and the hyperparameters of the GP prior.
First, the choice of $\phi$ governs the trade-off between exploitation and exploration in GP-UJB. Computationally, it is advantageous to select $\phi$ that yields a closed-form expression of GP-UJB. Table˜1 summarizes examples of such choices. Modeling the log unnormalized density $\log q(\theta)$ via GPs is a well-established approach in the literature (e.g., Osborne2012; Jarvenpaa2021). Following these precedents, we adopt the log transformation $\phi^{-1}(\cdot)=\log(\cdot)$ for the inverse, meaning that $\phi(\cdot)=\exp(\cdot)$, as the default setting. We explore alternative choices in Section˜C.2.
Table 1: Closed-form expressions of GP-UJB for selected functions $\phi$. Here, $\varphi$ and $\Phi$ denote the probability density function and the cumulative distribution function of the standard normal distribution, respectively, and we define $\sigma_{n}(\theta):=\sqrt{k_{n}(\theta,\theta)}$ for better presentation.
@@
$\max(0,x)$ $x$ $m_{n}(\theta)\Phi\left(m_{n}(\theta)/\sigma_{n}(\theta)\right)+\sigma_{n}(\theta)\varphi\left(m_{n}(\theta)/\sigma_{n}(\theta)\right)$
@@
Next, the mean and covariance functions of the GP prior typically depend on a set of hyperparameters $\psi$, such as the length scale and variance of a Gaussian kernel. Properly tuning $\psi$ is critical for the performance of GP regression. In the context of BO, updating $\psi$ via maximum likelihood estimation at each iteration is standard practice to maximize the performance. Accordingly, we update $\psi$ for GP-UJB by maximizing the log-likelihood of the GP posterior at each step.
@@
In sequential design settings, GP surrogate models may suffer from instability during the initial iterations due to overfitting to scarce data. To mitigate this, it is common to select the first few points at random rather than optimization. Algorithm 1 readily accommodates this strategy: for a fixed integer $N_{0}>0$, the initial $N_{0}$ samples $\{\theta_{n}^{*}\}_{n=1}^{N_{0}}$ can be selected from the proposal sequence, such as the scaled Halton sequence, without optimization.
@@ 6 Empirical Assessment
This section examines the empirical performance of BIS across four distinct experiments, ranging from numerical validation to practical application. First, we begin with three benchmarks to validate sample quality and convergence of BIS. Next, we consider a weather forecast model from Lorenz1995, where Bayesian inference is performed via synthetic likelihoods based on simulations (Hakkarainen2012). We then apply BIS to exact Bayesian inference for g-and-k models (Prangle2020), a setting where density evaluation requires numerical optimization at each data point. Finally, we illustrate the practical utility of our approach through a real-world application, modeling precipitation data (Kaufman2008) via an MRF. Source code for the experiments is available at https://github.com/takuomatsubara/Bandit-Importance-Sampling.
@@ 6.1 Benchmark Densities
We evaluate BIS on three bivariate benchmark densities, each designed to pose a distinct challenge. The first is a simple unimodal Gaussian density. The second is a bimodal density, representing a scenario with multiple modes that requires efficient domain exploration. The third is a banana-shaped density, illustrating a case with complex tail-probability geometries. All the three densities are defined in the following form proposed in Jarvenpaa2021:
[…]
where the definitions of $T_{1}$, $T_{2}$, and $\rho$ for each density are provided in Section˜D.1. The domains are defined as $\Theta=[-16,16]^{2}$, $\Theta=[-6,6]^{2}$, and $\Theta=[-6,6]\times[-20,2]$, respectively. The aim of this experiment is to demonstrate the performance of BIS under each benchmark scenario.
@@
Figure 3: Visualization of 100 samples generated by BIS for each density. The top panels show the contour colourmap of each density. The bottom panels show the samples, distinguishing between the initial 10 points (stars) and the subsequent 90 points selected by BIS (circles). Solid lines indicate density contours, where the values are raised to the power of $1/3$ to accentuate the tail geometry.
For each target density, we generated 100 weighted samples using BIS. The initial 10 points were selected from a Halton sequence to ensure the stability of the GP surrogate model, following the protocol in Jarvenpaa2021. The GP-UJB prior employed a zero mean function and a Gaussian kernel, treating the length scale and variance as tunable hyperparameters. The candidate pool size was fixed at $M=2048$, where a sensitivity analysis regarding this parameter is provided in Section˜C.3. Figure˜3 presents a comparison between the target densities and the obtained samples, demonstrating that BIS effectively identifies the high-probability regions. Notably, the method successfully resolved the complex geometries of the bimodal and banana-shaped densities.
We quantified the approximation quality of the obtained samples using the MMD. Figure˜4 displays the MMD between the target density and the BIS samples. For comparison, we evaluated a standard importance sampling baseline using the scaled Halton sequence with self-normalized weights. Section˜D.1 investigated the number of samples required for the standard importance sampling baseline to match the approximation error of 100 BIS samples. On average, BIS reduces the sample size required to achieve a comparable error by nearly 95%.
@@
Figure 4: Approximation error for BIS (solid line) and standard importance sampling (dotted line). Shaded regions indicate 95% confidence intervals over 10 independent runs.
@@
Figure 5: Approximation error of GP-based density estimators trained on input points selected by BIS (solid line), QMC (dotted line), the randomized BO (dashed line), and the EIV design (dash-dotted line). Shaded regions indicate 95% confidence intervals over 10 independent runs.
BIS does not require the GP surrogate to converge to the target density, where a reasonable approximation suffices for guiding sample selection. Nonetheless, assessing the GP surrogate’s fidelity provides valuable insight. To this end, we evaluated the total variation distance (TVD) between the target density and the density estimator derived by plugging the GP mean $m_{n}(\theta)$ in the exponential map. Figure˜5 illustrates the decay of the approximation error. We compared the estimator trained on the BIS samples against the estimator trained on three alternative designs. The first is standard QMC samples. The second is the expected integrated variance (EIV) design of Jarvenpaa2021, which modifies Sinsbeck2017 for better numerical stability. The last is the randomized BO of Kim2024. The distinction lies solely in the selection of input points; the GP hyperparameters were identical across all methods. As demonstrated in Figure˜5, the GP surrogate trained on BIS samples consistently yields the lowest TVD across all cases.